# RaanA: A Fast, Flexible, and Data-Efficient Post-Training Quantization Algorithm

### v1

@@ 1 Introduction
    Strong Performance: RaanA performs comparably to state-of-the-art PTQ methods. Notably, it remains effective even at extremely low bit-widths (e.g., $<3$ average bits), a regime where many other methods struggle.
@@ 2 Background and Related Work
There have been a vast literature of vector quantization in ANNS. The classical scalar quantization (SQ) and its variant LVQ (Aguerrebere et al., 2023) take finite uniform grids as codebooks and independently round every coordinate to an unsigned integer. These methods support to compute inner product based on arithmetics of unsigned integers without decompression. However, they failed to achieve reasonable accuracy using low-bit quantization ($<4$ bits). PQ (Jegou et al., 2011) and OPQ (Ge et al., 2013) construct a codebook via KMeans clustering and quantizes vectors by mapping them to the nearest vector in the codebook. With clustering, these methods better capture the distributions of a dataset and provide better accuracy using low-bit quantization. However, their computation of inner product replies on looking up tables, which is costly in modern systems. RaBitQ (Gao & Long, 2024) and its multi-bit extension (Gao et al., 2024) construct a codebook by translating, normalizing, and randomly rotating finite uniform grids. They achieve superior accuracy while enabling efficient computation. Theoretically, they provide an unbiased estimator with asymptotically optimal error bounds for the space-accuracy trade-off in inner product estimation between unit vectors. See Section A.2 for more details.
@@ 3 Preliminaries
In this paper, we fix a trained neural network $f:\mathbb{R}^{n\times d_{0}}\times\mathbb{R}^{m}\to\mathbb{R}$ as the object of study, where $d_{0}$ is the dimensionality of the input and $m$ is the number of learnable parameters [^1] A neural network with multi-dimensional output can be viewed as several neural networks with scalar output, and therefore we only consider NNs with scalar output.. We use ${\boldsymbol{x}}\in\mathbb{R}^{n\times d_{0}}$ and ${\boldsymbol{\theta}}$ to denote the input data and the collection of all parameters of $f$, respectively, where $n$ is the token $n$ can also be understood as the batch size, since we only consider linear layers that do not involve any cross-token interaction..
@@ 3.1 Overall Framework
Algorithm 1: The overall framework of RaanA. $\mathsf{Err}_{k}(b;{\boldsymbol{x}})$ is an estimation of the overall error brought by the $k$-th linear layer with $b$-bit quantization, estimated at point ${\boldsymbol{x}}$, and RaBitQ-H is the RaBitQ-H quantization algorithm. $m_{k}=d_{k}c_{k}$ is the number of parameters at layer $k$; $n_{c}$ is the number of calibration data; $\mathscr{B}$ is a set containing all desired choice of layer-wise bit-width (e.g. =$\mathscr{B}=\{1,2,\cdots 8\}$), and $R$ is the desired total number of bits used (i.e. bits per parameter times total number of weight parameters).
@@ 5 RaBitQ-H: A variant of RaBitQ with Randomized Hadamard Transformation
${\boldsymbol{D}}^{(k)}\leftarrow\mathop{\mathrm{diag}}\left({\boldsymbol{\xi}}\right)$;
@@
Remarks on Implementation. The practical implementation of RaBitQ-H is slightly more complicated than described here, involving two additional components. First, since the fast Hadamard transformation only applies to vector sizes that are powers of 2, we need to handle vectors whose sizes are not powers of 2. Second, we use small implementation tricks (such as input centralization) that preprocess the inputs and weights before running RaBitQ-H; these tricks do not affect the output of the quantization / de-quantization algorithm but in practice reduce the error rate. Due to space limitations, we defer these details to Sections C.2 and C.3.
@@ 6 Experiment Results
In this section, we present our experimental results. We focus on the performance loss after quantization, compared to the full-precision (fp16) model. Following previous work (Lin et al., 2024; Shao et al., 2023), we evaluate our method on the LLaMA model family and language modeling tasks.
Datasets.
Baseline Methods.
We compare RaanA with GPTQ (Frantar et al., 2023), AWQ (Lin et al., 2024), OmniQuant (Shao et al., 2023), and Quip#no FT & no ${}_{\text{no FT \& no $E_{8}$}}$ (Tseng et al., 2024) [^5] The full version of Quip# (Tseng et al., 2024) fine-tunes the model after quantization. In this paper, we compare RaanA with Quip#no FT & no ${}_{\text{no FT \& no $E_{8}$}}$, the version without fine-tuning, since fine-tuning is a universal plug-in component in quantization and is orthogonal to our contribution.. Additionally, for 4-bit quantization, we also compare with EasyQuant (Tang et al., 2024), a lightweight quantization method that does not require calibration data and targets 4-bit quantization. Note that most baseline models use tricks such as grouping and keeping full-precision outliers that introduce extra bit costs [^6] The only exception in our baselines is Quip#no FT & no ${}_{\text{no FT \& no $E_{8}$}}$, which has a precise control over the average number of bits. We admittedly require slightly more bits to match its performance. However, RaanA is significantly more lightweight than Quip#no FT & no ${}_{\text{no FT \& no $E_{8}$}}$, which uses over 6000 calibration samples to estimate the layer-wise Hessian., generally ranging from 0.1 to 0.3 bits. To make a fair comparison, we report RaanA performance with $x+0.1$ bits and $x+0.3$ bits for $x\in{2,3,4}$.
6.1 Performance
@@
Table 1: Perplexity results on wikitext2. Methods labeled ‘+” in the number of bits use tricks such as grouping and keeping outliers full-precision that bring $0.1\sim 0.3$ extra bit cost. Quip# refers to Quip#no FT & no ${}_{\text{no FT \& no $E_{8}$}}$. For OmniQuant, GPTQ and AWQ, we compare with the version with grouping size $128$. Best performance of each category is labeled bold.
@@ 6.3 Quantization Time
We report the time RaanA used for quantization in Table 3 under a specific setting as an illustration of the efficiency of RaanA. The experiments are conducted with 4 NVIDIA A100 GPUs [^7] Here the GPU configuration does not significantly impact the quantization time since the bottleneck is the CPU computation of RaBitQ. The machine we use has two AMD EPYC 7513 CPUs.. As shown, RaanA completes the quantization of a 70B model in under one hour, significantly faster than other heavyweight quantization methods such as Quip#no FT & no ${}_{\text{no FT \& no $E_{8}$}}$, which can take up to 10 hours for the same model size, despite having comparable performance.
@@ 7 Discussion and Conclusion
Last but not least, we would like to advocate future large language model designers to use a powers of 2 as the hidden size more often; as arbitrary as it may seem, this choice actually has the following advantages that make quantization easier and faster: 1) It maximizes the GCD between layers and thus improves the speed of the AllocateBits algorithm. 2) It makes it easier to use fast Hadamard Transformations since it is only defined for spaces whose dimension is a power of $2$.
@@ A.2 Detailed Introduction to RaBitQ
In this sub-section, we provide a brief introduction to RaBitQ (Gao & Long, 2024) and its multi-bit extension (Gao et al., 2024). The RaBitQ methods were proposed initially for vector quantization in database systems. They target to produce accurate estimation of inner product and Euclidean distances based on the quantized vectors while using the minimum space for storing quantization codes. Specifically, let ${\boldsymbol{P}}$ be a Johnson-Lindenstrauss Transformation (a.k.a., random rotation) (Johnson & Lindenstrauss, 1984). For a vector ${\boldsymbol{x}}\in\mathbb{R}^{d}$, RaBitQ randomly rotates it into ${\boldsymbol{Px}}$ and quantizes ${\boldsymbol{Px}}$ to a vector of $b$-bit unsigned integers ${\boldsymbol{\bar{x}}}\in[2^{b}]^{d}$ with a rescaling factor $t\in\mathbb{R}$. Then for another vector ${\boldsymbol{y}}\in\mathbb{R}^{d}$, RaBitQ estimates the inner product ${\boldsymbol{x}},{\boldsymbol{y}}\right>$ with
[…]
@@
[…]
@@
[…]
@@ Appendix B Theoretical Results
$\left|g({\boldsymbol{h}})-g({\boldsymbol{h}}+{\boldsymbol{\epsilon}})\right|$ $\nabla g({\boldsymbol{h}}),{\boldsymbol{\epsilon}}\right>+O\left(\left\|\nabla^{2}g({\boldsymbol{h}})\right\|\|{\boldsymbol{\epsilon}}\|^{2}\right)$ (20)
@@ C.2 Randomized Hadamard Transformation for Arbitrary Dimensionality
As we mentioned in Section A.1, the fast Hadamard Transformation is only defined with vector dimensionality $d$ that is a power of $2$. However, in practice, it is not always satisfied. In previous work such as (Tseng et al., 2024), this issue is solved by finding the largest factor of $d$ which is a power of $2$, say $\tilde{d}$, and applying Hadamard Transformation block-wisely, with each block has size $\tilde{d}$. However, in our experiment, we found this method extremely inefficient. For example, for LLaMA models, there can be $>20$ blocks.
@@
       ${\boldsymbol{\xi}}^{(j)}\leftarrow\left\{\xi^{(j)}_{i}\right\}_{i=1}^{\hat{d}}$, where $\xi_{i}^{(j)}$ is sampled i.i.d. from the Rademacher distribution;
@@ C.3 Tricks used in Quantization
Practically we find these tricks perform differently across different settings. To keep the configuration consistent and avoid heavy hyper-parameter tuning, in all the experiments presented in this paper, we use Centralization and Column Outlier Excluding.
@@ Appendix D Additional Experiment Results
In this section we display additional experiment results. Table 4 contains perplexity comparison between RaanA with baseline methods on c4, where RaanA uses few-shot calibration. Table 5 contains the few-shot vs zero-shot comparison of RaanA on c4. These additional experimental results support our claim in main paper.

### v2

@@ 1 Introduction
    Strong Performance: RaanA performs comparably to state-of-the-art PTQ methods. Notably, it remains effective even at extremely low bit-widths (e.g., <3<3 average bits), a regime where many other methods struggle.
@@ 2 Background and Related Work
There have been a vast literature of vector quantization in ANNS. The classical scalar quantization (SQ) and its variant LVQ (Aguerrebere et al., 2023) take finite uniform grids as codebooks and independently round every coordinate to an unsigned integer. These methods support to compute inner product based on arithmetics of unsigned integers without decompression. However, they failed to achieve reasonable accuracy using low-bit quantization (<4<4 bits). PQ (Jegou et al., 2011) and OPQ (Ge et al., 2013) construct a codebook via KMeans clustering and quantizes vectors by mapping them to the nearest vector in the codebook. With clustering, these methods better capture the distributions of a dataset and provide better accuracy using low-bit quantization. However, their computation of inner product replies on looking up tables, which is costly in modern systems. RaBitQ (Gao & Long, 2024) and its multi-bit extension (Gao et al., 2024) construct a codebook by translating, normalizing, and randomly rotating finite uniform grids. They achieve superior accuracy while enabling efficient computation. Theoretically, they provide an unbiased estimator with asymptotically optimal error bounds for the space-accuracy trade-off in inner product estimation between unit vectors. See Section A.2 for more details.
@@ 3 Preliminaries
In this paper, we fix a trained neural network $f:\mathbb{R}^{n\times d_{0}}\times\mathbb{R}^{m}\to\mathbb{R}$ as the object of study, where $d_{0}$ is the dimensionality of the input and $m$ is the number of learnable parameters [^1] A neural network with multi-dimensional output can be viewed as several neural networks with scalar output, and therefore we only consider NNs with scalar output.. We use ${\boldsymbol{x}}\in\mathbb{R}^{n\times d_{0}}$ and $\boldsymbol{\theta}$} to denote the input data and the collection of all parameters of $f$, respectively, where $n$ is the token number222nn can also be understood as the batch size, since we only consider linear layers that do not involve any cross-token interaction..
@@ 3.1 Overall Framework
Algorithm 1: The overall framework of RaanA. $\mathsf{Err}_{k}(b;{\boldsymbol{x}})$ is an estimation of the overall error brought by the $k$-th linear layer with $b$-bit quantization, estimated at point $\boldsymbol{x}$}, and RaBitQ-H is the RaBitQ-H quantization algorithm. $m_{k}=d_{k}c_{k}$ is the number of parameters at layer $k$; $n_{c}$ is the number of calibration data; $\mathscr{B}$ is a set containing all desired choice of layer-wise bit-width (e.g. $\mathscr{B}=\{1,2,\cdots 8\}$), and $R$ is the desired total number of bits used (i.e. bits per parameter times total number of weight parameters).
@@ 5 RaBitQ-H: A variant of RaBitQ with Randomized Hadamard Transformation
0.4em ${\boldsymbol{D}}^{(k)}\leftarrow\mathop{\mathrm{diag}}\left({\boldsymbol{\xi}}\right)$;
@@
Remarks on Implementation.
The practical implementation of RaBitQ-H is slightly more complicated than described here, involving two additional components. First, since the fast Hadamard transformation only applies to vector sizes that are powers of 2, we need to handle vectors whose sizes are not powers of 2. Second, we use small implementation tricks (such as input centralization) that preprocess the inputs and weights before running RaBitQ-H; these tricks do not affect the output of the quantization / de-quantization algorithm but in practice reduce the error rate. Due to space limitations, we defer these details to Sections C.2 and C.3.
@@ 6 Experiment Results
[…]
In this section, we present our experimental results. We focus on the performance loss after quantization, compared to the full-precision (fp16) model. Following previous work (Lin et al., 2024; Shao et al., 2023), we evaluate our method on the LLaMA model family (Touvron et al., 2023) and the Qwen3 (Yang et al., 2025) model with language modeling tasks. Due to limited space, we only show results on LLaMA models in the main paper and defer results with Qwen3 to the appendix.
Datasets.
Baseline Methods.
[…]
6.1 Performance
@@ 6.3 Quantization Time
[…]
6.4 Ablation Studies
In this subsection, we investigate the contribution of each component of RaanA by conducting experiments and analyses with individual components isolated. The results demonstrate that both RabitQ-H and AllocateBits play indispensable roles in the overall performance of RaanA.
Ablation with AllocateBits
To examine the effect of AllocateBits, we compare the performance of the full RaanA model with that of a variant using uniform bit allocation (i.e., assigning the same number of bits to each parameter). The results, presented in Table 5, clearly show that AllocateBits significantly enhances performance—particularly in low-bit quantization—even when calibrated using only a single, arbitrarily chosen sentence.
RabitQ RabitQ-H w/o quant
Extra Time $\Theta(nd^{2})$ $\Theta(nd\log d)$ $\Theta(ndc)$
Extra bpm $16d/c$ $1/c$ $16$
Table 4: De-quantization time and extra bpm requirement comparison between RabitQ and RabitQ-H. The “w/o quant” shows the original time and bpm required if without quantization, assuming the original model is using 16-bits float point number.
Bits Allocation Calibration 2 bits 3 bits 4 bits
uniform - 1253.78 7.63 5.8
AllocateBits few-shot 71.52 6.43 5.72
AllocateBits zero-shot 46.33 6.60 5.77
Table 5: Perplexity results under different bits allocation algorithms. Experiments are performed with llama2-7b on wikitext2. Notice that since the calibration is only needed in AllocateBits stage, no calibration is needed after it is replaced with uniform bits allocation.
Ablation with RabitQ-H
The RHT module in RabitQ-H primarily improves de-quantization efficiency by reducing both computation time and the number of additional bits required to store de-quantization information. To highlight this improvement, Table 5 reports the de-quantization time complexity and extra bits per parameter (bpm) needed for a linear layer $f(\mathbf{X})=\mathbf{X}\mathbf{W}$, where $\mathbf{X}$ has shape $n\times d$ and $\mathbf{W}$ has shape $d\times c$. We compare the overhead of RabitQ and RabitQ-H against the cost of matrix multiplication in the original (non-quantized) model. The analysis reveals that, with RabitQ-H, both the extra time and memory overhead introduced by quantization are negligible since they are on the order of $O(1/c)$ or $O(\log d/c)$ relative to the original model. In contrast, the original RabitQ introduces an overhead of $\Theta(d/c)$, which becomes prohibitively large in practical settings where $d\approx c$.
@@ 7 Discussion and Conclusion
Last but not least, we would like to advocate future large language model designers to use a powers of 2 as the hidden size more often; as arbitrary as it may seem, this choice actually has the following advantages that make quantization easier and faster: 1) It maximizes the GCD between layers and thus improves the speed of the AllocateBits algorithm. 2) It makes it easier to use fast Hadamard Transformations since fast Hadamard Transformation is only defined for spaces whose dimension is a power of $2$.
@@ A.2 Detailed Introduction to RaBitQ
In this sub-section, we provide a brief introduction to RaBitQ (Gao & Long, 2024) and its multi-bit extension (Gao et al., 2024). The RaBitQ methods were proposed initially for vector quantization in database systems. They target to produce accurate estimation of inner product and Euclidean distances based on the quantized vectors while using the minimum space for storing quantization codes. Specifically, let $\boldsymbol{P}$} be a Johnson-Lindenstrauss Transformation (a.k.a., random rotation) (Johnson & Lindenstrauss, 1984). For a vector ${\boldsymbol{x}}\in\mathbb{R}^{d}$, RaBitQ randomly rotates it into $\boldsymbol{Px}$} and quantizes $\boldsymbol{Px}$} to a vector of $b$-bit unsigned integers ${\boldsymbol{\bar{x}}}\in[2^{b}]^{d}$ with a rescaling factor $t\in\mathbb{R}$. Then for another vector ${\boldsymbol{y}}\in\mathbb{R}^{d}$, RaBitQ estimates the inner product $\left<{\boldsymbol{x}},{\boldsymbol{y}}\right$> with
$\left<{\boldsymbol{x}},{\boldsymbol{y}}\right>\approx\left<{t\cdot(\boldsymbol{\bar{x}}-c_{b}\cdot\mathbf{1}_{d})},{\boldsymbol{Py}}\right$> (9)
@@
$\big|\left<{\boldsymbol{x}},{\boldsymbol{y}}\right>-\left<{t(\boldsymbol{\bar{x}}-c_{b}\mathbf{1}_{d})},{\boldsymbol{Py}}\right>\big|<\epsilon\|{\boldsymbol{x}}\|\|{\boldsymbol{y}}\|$ (10)
@@
$\big|\left<{\boldsymbol{x}},{\boldsymbol{y}}\right>-\left<{t(\boldsymbol{\bar{x}}-c_{b}\mathbf{1}_{d})},{\boldsymbol{Py}}\right>\big|<\frac{c_{\text{error}}}{\sqrt{d}2^{b}}\|{\boldsymbol{x}}\|\|{\boldsymbol{y}}\|$ (11)
@@ Appendix B Theoretical Results
$\left|g({\boldsymbol{h}})-g({\boldsymbol{h}}+{\boldsymbol{\epsilon}})\right|$ $\leq\left<\nabla g({\boldsymbol{h}}),{\boldsymbol{\epsilon}}\right>+O\left(\left\|\nabla^{2}g({\boldsymbol{h}})\right\|\|{\boldsymbol{\epsilon}}\|^{2}\right)$ (20)
@@ C.2 Randomized Hadamard Transformation for Arbitrary Dimensionality
As we mentioned in Section A.1, the fast Hadamard Transformation is only defined with vector dimensionality $d$ that is a power of $2$. However, in practice, it is not always satisfied. In previous work such as (Tseng et al., 2024), this issue is solved by finding the largest factor of $d$ which is a power of $2$, say $\tilde{d}$, and applying Hadamard Transformation block-wisely, with each block has size $\tilde{d}$. However, in our experiment, we found this method extremely inefficient. For example, for LLaMA models, there can be >20>20 blocks.
@@
 $^{\boldsymbol{\xi}}^{(j)}\leftarrow\left\{\xi^{(j)}_{i}\right\}_{i=1}^{\hat{d}}$, where $\xi_{i}^{(j)}$ is sampled i.i.d. from the Rademacher distribution;
@@ C.3 Tricks used in Quantization
Practically we find these tricks have incremental contribution on the performance, and different combinations of them perform differently across settings. Below in Section C.4 we perform a preliminary experiment on their impact to the performance. In all the other experiments presented in this paper, to keep the configuration consistent and avoid heavy hyper-parameter tuning, we use Centralization and Column Outlier Excluding.
C.4 Ablation Studies on Implementation Tricks
In order to show the contribution of the implementation tricks to the performance, in this subsection we compare their impact to the performance with different combinations. The results are shown in Table 6. It can be easily inferred from the table that these implementation tricks only have incremental contribution to the overall performance.
Combination of tricks few (3.3bits) zero (3.3bits) few (4.3bits) zero (4.3bits)
(none) 11.35 12.20 10.35 10.49
cent 11.14 11.93 10.33 10.45
cent, col-out 11.05 11.56 10.06 10.27
cent, row-out 10.98 11.83 10.24 10.51
cent, col-out, row-out 11.02 11.49 10.14 10.25
col-out 11.18 11.71 10.21 10.32
row-out 11.24 11.93 10.23 10.37
col-out, row-out 11.13 11.63 10.24 10.32
Table 6: Perplexity comparison between different combinations of the implementation tricks. All experiments are performed with qwen-8b on wikitext2. In the table, “cent“ refers to Centralization, “row-out” refers to Row Outlier Excluding and “col-out” refers to Column Outlier Excluding.
@@ Appendix D Additional Experiment Results
In this section we present additional experiment results on extra models / datasets. These additional experimental results clearly support our claim in main paper.
Table 7 presents the perplexity results of RaanA with Qwen3 on wikitext2 dataset, compared with GPTQ and AWQ with group size 128. The baseline results are taken from (Zheng et al., 2025).
Method Avg. bits qwen-4b qwen-8b qwen-14b
fp16 16 13.7 9.71 8.64
AWQ 2+ 1.38E7 1.21E7 6.06E6
GPTQ 2+ 1.95E2 52.1 25.5
RaanA 2.3 328.20 38.13 17.48
AWQ 3+ 91.0 27.5 19.4
RaanA 3.1 20.40 12.05 10.00
RaanA 3.3 18.63 11.14 9.79
AWQ 4+ 18.1 11.3 9.48
GPTQ 4+ 13.5 9.96 8.88
RaanA 4.1 15.4 10.18 9.23
RaanA 4.3 14.9 9.97 9.18
Table 7: Perplexity results of Qwen3 on wikitext2. The setting and format are the same as Table 1, except model.
Table 8 contains perplexity comparison between RaanA with baseline methods with LLaMA models on c4, where RaanA uses few-shot calibration. Table 9 contains the few-shot vs zero-shot comparison of RaanA on c4.